Which of the following is the better buy?
a. 67 cents for 58 ounces or 17 cents for 15 ounces
b. 29 ounces for 13 cents or 56 ounces for 27 cents
c. 17 ounces for 23 cents, 25 ounces for 34 cents, or 73 ounces for 96 cents

Answer by Boreal(15235)

cents per ounce or ounces per cent.
1. 67/58=1.155; 17/15=1.133. The second is fewer cents per ounce. It is a better buy.
2. 29/13=2.31; 56/27=2.07. The first gives more ounces per cent. It is a better buy.
3. 17/23=0.74; 25/34=0.735; 73/96=0.760. The third is more ounces per cent. It is the best buy.